OpenStudy (anonymous):

What is the y-intercept of the equation? y = 2.1(0.15)x

OpenStudy (zrsmith):

The equation is y = mx +b, where b is the y-intercept. In this case, you have the m=slope=2.1(0.15), the variable x and nothing in the b (y-intercept). Therefore the y-intercept is 0. It might help to understand what the y-intercept is in this style of problem. The y-intercept is the place where the line crosses the y axis, and you can solve for this by making the x variable = 0. If you use the equation above, making your equation y = 2.1(0.15)x + b, and set x = 0 and solve for b, you will see that the solution is b = 0.

OpenStudy (unklerhaukus):

\[y(x) = 2.1\times(0.15)^x\] the y-intercept is when x=0 \[y(0) = 2.1\times(0.15)^0=\quad\dots\]